OpenStudy (jocelynstalnaker):

Will give medal or whatever it is you need! Its 6 questions long for one problem.

OpenStudy (anonymous):

What is it you need help in ?

OpenStudy (anonymous):

must be long.

OpenStudy (jocelynstalnaker):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. -You may graph your equation by hand on a piece of paper and scan your work. -You may graph your equation using graphic technology that can be found in the Course Information area. 5. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. 6. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (anonymous):

Wow o.o hmmm hold on 1 sec i might have an answer for you ^_^

OpenStudy (anonymous):

what do u understand by a linear equation? and what do you know about slopes?

OpenStudy (jocelynstalnaker):

Ok, Linear equation is a function that is defined by a linear equation, I believe and I dont really know a lot about slopes

OpenStudy (jocelynstalnaker):

A slope can either decrease or increase, I think.

OpenStudy (anonymous):

ok so umm you're familiar with "y=mx + c" ... right?

OpenStudy (jocelynstalnaker):

Yes

OpenStudy (jocelynstalnaker):

Wait, no its + b

OpenStudy (anonymous):

"Type the first equation from the group you selected..." what group?

OpenStudy (jocelynstalnaker):

1200=2x+3y

OpenStudy (jocelynstalnaker):

SO 0=2x+3y-1200 -3y=2x-1200 divide everything by (-3) y=(-2/3)x+400 m (slope)= -2/3 b (y intercept)= 400 Right?

OpenStudy (anonymous):

hmm y = negative three - halvesx +4 = -3 - (1/2)x + 4 y = -3 - (1/2)x + 4 y = -(1/2)x + 4 - 3 y = -(1/2)x + 1 do you agree?

OpenStudy (jocelynstalnaker):

Sorry it took so long, my laptop crashed. No.. I dont understand

OpenStudy (anonymous):

hmm okay, ill help, sorry i haven't replied takes forever for the screen to load.

OpenStudy (anonymous):

1. 2x + 3y = 1200 2x-1200=-3y 1/3*(2x-1200)=1/3*-3y 2/3x-400=-y -2/3x+400=y

OpenStudy (anonymous):

The second one you would need to find the slope intercept which was given above, -2/3... 2. I would put a dot on (0,400) then I would use the slope to do the rest or go 2 points to the right and 3 to the down. After this it should look like 3,398.

OpenStudy (jocelynstalnaker):

Dont worry its my computer too.

OpenStudy (anonymous):

3. F(x) = 2x + 1200

OpenStudy (anonymous):

Can you do the graph by yourself?

OpenStudy (jocelynstalnaker):

Yes, Thank you so much! Could you help me on number 5 and 6?

OpenStudy (anonymous):

5. 2x + 3y = 1500 -3y = 2x – 1500 1/3* -3y = (2x – 1500)*1/3 -2/3x + 500 = y

OpenStudy (anonymous):

6. F(x)=mx + b F(x) = mx + 300 F(x) = 6/9x + 300 F(x) = 2/3x + 300

OpenStudy (jocelynstalnaker):

Thank you so much!

OpenStudy (anonymous):

Please review this and learn :D and a pleasure to help you.

OpenStudy (jocelynstalnaker):

Actually i need help on 3): @ILuvHomework

OpenStudy (anonymous):

the graph is for 4, is that the one you need help with ?